A project has unused construction funds that can be deposited into an interest bearing account for later use. The annual effective interest rate of the account is 18%, and interest is compounded monthly. Approximately how much money must be deposited to yield $60,000 at the end of 11 months?

a. $48,750

b. $50,100

c. $50,950

d. $51,550
Answer #1

monthly interest rate = 18% / 12 = 1.5%

F= P *(1+i)^t

60000 = P *(1+0.015)^11

60000 = P *(1.015)^11

P = 60000 / 1.015 ^11

P = 50935.99 ~ 50950

Option C is correct answer